### v1

# Accelerating Targeted Hard-Label Adversarial Attacks in Low-Query Black-Box Settings

Abstract
Deep neural networks for image classification remain vulnerable to adversarial examples – small, imperceptible perturbations that induce misclassifications. In black-box settings, where only the final prediction is accessible, crafting targeted attacks that aim to misclassify into a specific target class is particularly challenging due to narrow decision regions. Current state-of-the-art methods often exploit the geometric properties of the decision boundary separating a source image and a target image rather than incorporating information from the images themselves. In contrast, we propose Targeted Edge-informed Attack (TEA), a novel attack that utilizes edge information from the target image to carefully perturb it, thereby producing an adversarial image that is closer to the source image while still achieving the desired target classification. Our approach consistently outperforms current state-of-the-art methods across different models in low query settings (nearly 70% fewer queries are used), a scenario especially relevant in real-world applications with limited queries and black-box access. Furthermore, by efficiently generating a suitable adversarial example, TEA provides an improved target initialization for established geometry-based attacks.
1 Introduction
Deep neural networks have achieved remarkable performance in image classification tasks, powering applications from autonomous systems [2, 5] to medical diagnostics [9, 12]. However, they have repeatedly been shown to be vulnerable to adversarial examples [11, 23]. These are small, often imperceptible perturbations to a correctly classified image that cause a misclassification. Although many of these attacks assume white-box access to a model’s internals, hard-label (decision-based) attacks offer a more challenging yet practical setting where only the top-1 predicted label is observed. This limited-feedback scenario commonly arises in commercial APIs [13]. In this realm, targeted attacks, which push the model’s prediction to a specific target class, are inherently more difficult since the decision regions corresponding to the specific target classes are usually narrower and more isolated.
In black-box settings, targeted hard-label attacks have become an active area of research, with several techniques proposed in the literature, with state-of-the-art methods relying on the geometry of the decision boundary separating the source image from the target class. Geometry-informed attacks traverse in a lower-dimensional space and fall into two categories: Boundary Tracing Attacks and Gradient Estimation Attacks. Boundary Tracing Attacks ([3, 4, 18, 22]) perform walks along the decision boundary while Gradient Estimation Attacks ([6, 14, 16, 17, 19, 24, 25]) perform the same walks but using information about the approximate tangent/normal to the decision boundary in a local neighborhood to their adversarial image. Although powerful for local refinement, both approaches tend to burn through queries when the source image lies far from a given adversarial image in a target class, wasting many queries before reaching a narrow region where local geometry can more effectively be leveraged.
Further, in practice, many real-world scenarios such as commercial pay-per-query APIs, impose severe constraints on the number of queries that can be made to a target model in a specified time. Often practical query limits may be on the order of a few hundred to fewer than a few thousand queries. Under these conditions, the limited feedback available makes it difficult to effectively use information about the local decision boundary geometry in the early stages of an attack. When the decision space is still wide, movement along restricted lower dimensional spaces leads to limited incremental progress, and gradient estimation methods often use queries in estimating local gradients by sampling predictions in a local neighborhood instead of moving. This raises a pressing need to develop methods that are efficient in the low-query regime. In a high query setting, this would also lead us to a good starting point to employ the geometry-informed methods since we arrive at a good adversarial point quickly and can use the geometry-informed information more effectively. To this end, we propose Targeted Edge-Informed Attack (TEA), a novel targeted adversarial attack designed for hard-label black-box scenarios within a restricted query budget. Rather than depending on the local properties of the decision boundary, TEA leverages the intrinsic features of the target image itself - specifically, the edge information obtained using Sobel filters [21] help identify prominent structural features in the target image. Edges encode high-magnitude spatial gradients that delineate object boundaries [21]. Research has shown that early layers fire on oriented edge filters [1, 26], and that shape/edge cues remain predictive even when textures are suppressed [10]. The core idea is to preserve these low-level features, while applying perturbations to the non-edge regions of an image, allowing us to stay in the target class while pushing the adversarial image towards the source image. When our progress plateaus, evidenced by a series of consecutive queries that fail to achieve further reduction in distance while maintaining target class prediction, one can switch to current state-of-the-art geometry-informed methods for a local refinement procedure. Hence we make the following contributions:
Global initialization for low-query efficiency: We introduce an edge-informed perturbation strategy, enabling rapid progress toward the source image in the early stages of an adversarial attack. This strategy involves a two-step process: First, a global edge-informed search is performed, and then, edge-informed updates are applied to small patches using Gaussian weights.
Empirical validation under strict query budgets: We perform extensive evaluations on the ImageNet validation dataset [7] across four architectures (ResNet-50 [12], ResNet-101 [12], VGG16 [20], and ViT [8]). Our attack consistently outperforms existing state-of-the-art hard-label methods, including HSJA, Adaptive History-driven Attack (AHA) [15], CGBA, and CGBA-H, under realistic query budgets (fewer than 1000 queries). To achieve a 60% reduction in distance from a target image to a source image, TEA required on average 251 queries across the four models - 70% fewer than AHA (the second fastest), which required 845 queries.
The rest of this paper is structured as follows. Section 2 reviews related targeted hard-label adversarial attacks. Section 3 describes our methodology, while Section 4 presents experimental results. Finally, Section 5 discusses potential limitations and outlines directions for future research.
2 Related Works
Early work in the realm of targeted hard-label adversarial attacks consisted of seminal work: Boundary Attack (BA) [3], which proposed a method of traversing the decision boundary that separates an adversarial image from a source image. Building on this framework, BiasedBA [4] incorporated directional priors, such as perceptual and low-frequency biases, to restrict the search to more promising regions. BA with Output Diversification Strategy (BAODS) [24] integrates diverse gradient-like signals into the exploration process of BA, thereby strengthening the original method.
Following these foundational methods, Hybrid Attack (HA) [22] employed a combination of heuristic search strategies to balance global exploration and local refinement. Advancing toward more gradient-centric techniques, qFool [16] leverages the local flatness of decision boundaries to streamline the attack process. Meanwhile, HopSkipJumpAttack (HSJA) [6] locally approximates the normal to the decision boundary to “jump off” from it before progressing toward the source image. Complementarily, Tangent Attack exploits locally estimated tangents of the decision surface to steer the adversarial perturbation toward the source image. Addressing the challenges posed by high-dimensional input spaces, Query Efficient Boundary Attack (QEBA) [14] projects gradient estimation into lower-dimensional subspaces, such as the frequency domain, thus significantly reducing the number of required queries.
Incorporating historical query information, Adaptive History-driven Attack (AHA) [15] adapts its search trajectory based on previous successes and failures, while Decision-based query Efficient Adversarial Attack based on boundary Learning (DEAL) [6] employs an evolutionary strategy that concentrates queries on promising regions of the input space. SurFree [18] demonstrated that using 2D planes and semicircular trajectories toward the source image was an effective strategy. This was based on the fact that under the assumption of a flat decision boundary, the point on the boundary that is closest to the source image is precisely where the semicircular trajectory intersects it. Building on this idea, Curvature-Aware Geometric black-box Attack (CGBA) and its variant that is more suited for targeted attacks - CGBA-H [19], use normal estimation at the decision boundary to select the traversal plane. To the best of our knowledge, CGBA, and CGBA-H serve as the current state of the art in decision-based adversarial attacks.
3 Methodology
In this section, we formalize the hard-label attack setting and detail our attack, which drastically reduces early query cost, especially when the target image is far from the source image, and exists in a wider decision space. Once the adversarial image achieves a good distance and reaches a narrow decision space, one can use the image as an initialization and continue refining with existing geometric-based methods.
Problem Statement. We consider a hard-label image classifier
$f:[0,1]^{C\times H\times W}\to\mathbb{R}^{K},$ (1)
where $C$ denotes the number of color channels, $H$ and $W$ are the image height and width, and the classifier distinguishes among $K$ classes. For any query image $x$, we do not observe its continuous output (e.g., logits or probabilities) but only the predicted label index
$\hat{y}(x)=\arg\max_{1\leq k\leq K}[f(x)]_{k}.$ (2)
Let $x_{s}$ be a source image correctly classified as $y_{s}$. In a targeted attack, we start with a target image $x_{t}$ (correctly classified as $y_{t}$). Our goal is to find an adversarial image $x_{\mathrm{adv}}$ that is as close as possible to $x_{s}$ (in the $\ell_{2}$ norm) while maintaining the target classification, i.e.,
$x_{\mathrm{adv}}=\arg\min_{x}\|x-x_{s}\|_{2},\quad\text{subject to}\quad\hat{y}(x)=y_{t}.$ (3)
We propose a two-part procedure for perturbing the target image $x_{t}$ to get closer to $x_{s}$ while maintaining adversarial requirements. First, a global edge-informed search coarsely aligns major image regions. Second, a patch-based edge-informed search further perturbs local regions in our adversarial image. Note that neither step leverages local decision boundary geometry or gradient estimation, which are typically beneficial only when the adversarial image lies near a narrow decision space and when movements towards the source image are unlikely to maintain classification. This approach avoids burning queries in a wider decision space.
3.1 Global Edge-Informed Search
The global perturbation step aims to coarsely align $x_{t}$ towards $x_{s}$ by modifying predominantly smooth, non-edge areas, thereby preserving crucial edge structures that help maintain target classification.
Soft Edge Mask. To this end, we detect edges in $x_{t}$ using the Sobel operator [21], as depicted in Figure 1. A subsequent blurring operation yields a soft edge mask $M_{\mathrm{edge}}$, where $M_{\mathrm{edge}}(i,j)\in[0,1]$ has values close to 1 at edge locations and gradually transitions to 0 in smoother regions. We describe this in Algorithm 1.
Algorithm 1 Soft Edge Mask
1:  Inputs: Image $x$, edge thresholds $\{T_{\ell},T_{h}\}$, Gaussian blur kernel size $b$, intensity factor $\gamma$, small constant $\epsilon$
2:  Output: Soft edge mask $M_{\mathrm{edge}}$
3:  $x^{\text{gray}}\leftarrow\text{GrayScale}(x)$
4:  $(s_{x},s_{y})\leftarrow(\text{Sobel}(x^{\text{gray}},0),\;\text{Sobel}(x^{\text{gray}},1))$
5:  $G\leftarrow(s_{x}^{2}+s_{y}^{2})^{1/2}$
6:  $G\leftarrow 255\cdot\left(G/(max_{i,j}(G)+\epsilon)\right)$
[…]
10:  $M_{\mathrm{edge}}\leftarrow\gamma\cdot\text{norm}$
11:  return  $M_{\mathrm{edge}}$
Global Interpolation. Starting from $x_{0}=x_{t}$, we perform iterative updates:
$x_{k+1}\leftarrow x_{k}+\alpha(x_{s}-x_{k})\odot\left(I-M_{\mathrm{edge}}\right),$ (4)
where $\odot$ denotes the element-wise (Hadamard) product. The scaling factor $\alpha$ is optimized via a momentum based search. Masking out the edge regions in this update ensures that the interpolation primarily affects smooth areas, thus preserving the overall structure necessary for target classification. The complete procedure is summarized in Algorithm 2. Once the improvements stagnate, we transition to a local, patch-based refinement.
Algorithm 2 Global Edge-Informed Search
1:  Inputs: Source image $x_{s}$, target image $x_{t}$, soft edge mask $M_{\mathrm{edge}}$, tolerance $\tau$, maximum queries $qc_{\max}$, initial step factor $\eta$, momentum $\mu$
2:  Output: Adversarial image $x_{\mathrm{adv}}$ such that $\hat{y}(x_{\mathrm{adv}})=y_{t}$
3:  $x_{\text{current}}\leftarrow x_{t}$, $v\leftarrow 0$, and $d\leftarrow x_{s}-x_{t}$
4:  Set step size: $s\leftarrow\|d\|_{2}\cdot\eta$, and initialize query count $qc\leftarrow 0$
5:  while $qc<qc_{\max}$ and $\|s\|\geq\tau$
6:     $v\leftarrow\mu\cdot v+(1-\mu)\cdot d$
7:     $x_{\text{next}}\leftarrow x_{\text{current}}+s\cdot(v\odot(I-M_{\mathrm{edge}}))$
8:     $qc\leftarrow qc+1$
9:      if $\hat{y}(x_{\text{next}})=y_{t}$
10:         $x_{\text{current}}\leftarrow x_{\text{next}}$
11:         $s\leftarrow 1.1\cdot s$
12:     else
13:         break
14:  return  $x_{\text{current}}$
3.2 Patch-Based Edge-Informed Search
After the global interpolation, some subregions of the image may still display significant discrepancies from $x_{s}$. In the patch-based refinement, we partition the image into patches
$\mathcal{P}\subseteq\{1,\dots,H\}\times\{1,\dots,W\},$ (5)
and for each patch, construct a local soft edge mask $M_{\mathcal{P}}$ analogously to $M_{\mathrm{edge}}$, but restricted to $\mathcal{P}$. Then, we apply a local interpolation:
$\widetilde{x}(\beta)=x_{k}+\beta(x_{s}-x_{k})\odot G\odot(I-M_{\mathcal{P}}),$ (6)
and search for the largest $\beta$ such that $\hat{y}(\widetilde{x}(\beta))=y_{t}$. Here, $G$ is a Gaussian weight applied across the patch to avoid artificial edges being created by our patches. This localized update is performed over multiple patches of varying sizes to steadily decrease the distance to the source image. The patch-based refinement is repeated until a termination criterion is met (e.g., 25 consecutive iterations with no further improvement). The corresponding illustration is depicted in Figure 1 and the relevant pseudocode is provided in Algorithm 3.
Figure 1: Overview of Patch-Based Edge-Informed Search. Edge information from the target image, obtained via the Sobel operator, is first blurred to generate a soft edge mask. A square patch is then selected and a Gaussian weighting function is applied. In the bottom right panel, the intensity of the modification is illustrated: dark red regions remain largely unchanged, while light green regions receive a more pronounced update. The lack of changes near the patch borders helps prevent the introduction of artificial edges.
Figure 2 offers a visual summary of TEA. The figure depicts the adversarial image throughout our method, and displays also a hotspot of individual pixel differences from the source image.
Figure 2: Visualization of TEA on a source–target image pair. The target image (initially classified as Bee-eater) is perturbed to resemble the source image (classified as Spoonbill), while preserving its original Bee-eater label. Global Edge-Informed Search efficiently applies edge-aware perturbations using only $15$ queries to achieve a $\approx$20reduction in distance to the source. Patch-Based Edge-Informed Search introduces localized, edge-aware modifications to small image regions, as seen in the hotspot of changes. Further refinement utilizing CGBA-H is illustrated in the narrow decision space.
Algorithm 3 Patch-Based Edge-Informed Search
1:  Inputs: Source image $x_{s}$, current adversarial image $x_{\mathrm{adv}}$, soft edge mask $M_{\mathrm{edge}}$ (from create_soft_edge_mask), minimum patch size $p_{\min}$, maximum patch size $p_{\max}$, step factor $\eta$, momentum $\mu$, maximum calls $N_{\max}$
2:  Output: Refined adversarial image $x_{\mathrm{adv}}$
3:  $n_{\text{break}}\leftarrow 0$
4:  while $n_{\text{break}}<25$
5:   $D\leftarrow\text{AvgPool}\left(|x_{s}-x_{\mathrm{adv}}|\right)$
6:   Select high-difference indices from $D$ and choose a random center $(i_{c},j_{c})$
7:   $p\leftarrow\mathrm{randInt}(p_{\min},\,p_{\max})$; $\mathcal{P}\leftarrow\{(i,j)\mid|i-i_{c}|\leq\lfloor p/2\rfloor,\;|j-j_{c}|\leq\lfloor p/2\rfloor\}$
8:   $M_{\mathcal{P}}(i,j)\leftarrow\begin{cases}1,&(i,j)\in\mathcal{P},\\ 0,&\text{otherwise.}\end{cases}$
9:   $x_{\text{patch}}\leftarrow\mathrm{Patch}(x_{\mathrm{adv}},\mathcal{P})$
10:   $m_{\text{patch}}\leftarrow 0,\quad d_{\text{base}}\leftarrow\|x_{s}-x_{\mathrm{adv}}\|_{2}$
11:   for iteration $=1$ to $N_{\max}$
12:    $d_{\text{local}}\leftarrow\mathrm{Patch}(x_{s},\mathcal{P})-x_{\text{patch}}$
13:    $m_{\text{patch}}\leftarrow\mu\cdot m_{\text{patch}}+(1-\mu)\cdot d_{\text{local}}$
14:    $s_{\text{patch}}\leftarrow\eta\cdot\|x_{s}-x_{\mathrm{adv}}\|_{2}$
15:    $G\leftarrow\mathrm{GaussianWeight}((i_{c},j_{c}),\sigma=p/3)$
16:    $\Delta x\leftarrow s_{\text{patch}}\cdot m_{\text{patch}}\odot\left(G\odot\left(1-(M_{\mathrm{edge}}\odot M_{\mathcal{P}})\right)\right)$
17:    $x_{\text{patch}}^{+}\leftarrow x_{\text{patch}}+\Delta x$
18:    $x_{\mathrm{temp}}\leftarrow x_{\mathrm{adv}}-x_{\mathrm{adv}}\odot M_{\mathcal{P}}+x_{\text{patch}}^{+}\odot M_{\mathcal{P}}$
19:    $d_{\text{new}}\leftarrow\|x_{s}-x_{\mathrm{temp}}\|_{2}$
20:    if $d_{\text{new}}\geq 0.999\cdot d_{\text{base}}$
21:     break
22:    if $\hat{y}(x_{\mathrm{temp}})=y_{t}$
23:     $x_{\text{patch}}\leftarrow x_{\text{patch}}^{+}$
24:     $x_{\mathrm{adv}}\leftarrow\mathrm{Replace}(x_{\mathrm{adv}},\mathcal{P},x_{\text{patch}})$
25:     $d_{\text{base}}\leftarrow d_{\text{new}}$
26:     $n_{\text{break}}\leftarrow 0$
27:    else
28:     $n_{\text{break}}\leftarrow n_{\text{break}}+1$
29:     break
30:  return $x_{\mathrm{adv}}$
4 Experiments
In this section, we present our empirical evaluation on the ImageNet ILSVRC2012 dataset, benchmarking TEA against existing targeted hard-label attack strategies. In what follows, we describe our experimental setup and metrics, then detail the performance of each method under a range of query budgets. Our findings indicate that TEA achieves efficient distance reduction to the source image, particularly in the early query regime, before switching to the current state-of-the-art geometry-based attack CGBA-H for further refinement in a narrow decision space.
4.1 Setup and Metrics
Computational Resources. The experiments were executed on a High-performance computing (HPC) Cluster. Each node on the cluster consisted of four NVIDIA GeForce GTX 1080 Ti GPUs (one GPU was allocated per source-target pair for a given attack).
Dataset and Image Pairs. We sample 1000 source-target image pairs from the ImageNet ILSVRC2012 validation set, ensuring the two images in each pair come from distinct classes. All images are resized to $3\times 224\times 224$. Each pair $(x_{s},x_{t})$ contains a source image $x_{s}$, correctly classified under its label, and a target image $x_{t}$, also correctly classified under a different label. The attack goal is to modify $x_{t}$ to approach $x_{s}$ under the $\ell_{2}$ norm while keeping the prediction unchanged.
Target Models. We evaluate our approach on four well-known classifiers: ResNet50 and ResNet101 (CNNs of varying depth), VGG16 (a CNN composed of repeated convolutional blocks), and ViT (a vision transformer that processes images as sequences of patches). These models represent a diverse set of architectures.
Compared Methods. We compare our method to four targeted hard-label attacks: HSJA, AHA, CGBA, and CGBA-H.For each method, the $\ell_{2}$ distance to the source image is recorded after each set of queries, and along with the classification label of the perturbed target image.
Evaluation Metrics. Performance is quantified using three metrics. First, we compute the $\ell_{2}$ distance from $x_{s}$ to the adversarial example generated for each of the 1000 image pairs as queries progress. Second, the attack success rate (ASR) is defined as the fraction of image pairs for which an $\alpha\%$ reduction in $\|x_{s}-x_{t}\|$ is achieved within a given number of queries. Third, we also measure the performance of each method when there is a set fixed low-query budget by measuring the ASRs for all $\alpha$ when each method has used up $500$ queries. Finally, we integrate the $\ell^{2}$ distance versus queries curve to obtain the area under the curve (AUC), which provides an aggregated measure of how rapidly the distance is reduced.
Implementation Details. Our implementation employs a two-stage process. In the first stage, we perform our proposed methodology TEA, where edge-aware distortions are applied to the target image while preserving its classification, allowing a rapid reduction in distance to the source image. In the second stage, once the perturbation has brought the image into a narrow decision space, the method switches to CGBA-H for further refinement. The last query performed with TEA is referred to as the turning point throughout the study. We switch to CGBA-H since it consistently performs better than other methods in a high query setting across different architectures.
4.2 Results
Average $\ell_{2}$-Distance vs. Queries. Table 1 presents a comparative analysis of the median $\ell_{2}$ distances achieved in the low-query regime—specifically, within the range defined by the average turning point computed over one thousand image pairs. This analysis underscores the rapid decrease in perturbation norm facilitated by TEA during the early query stages. Figure 3 extends this evaluation by depicting the average percentage reduction in the $\ell_{2}$ distance against the queries used.
Table 1: Median $\ell_{2}$ distances computed until the turning point across different architectures. Lower $\ell^{2}$ values denote faster adversarial example generation.
Query ResNet50 Query ResNet101
HSJA CGBA CGBA-H AHA TEA HSJA CGBA CGBA-H AHA TEA
100 92.797 88.857 83.104 80.438 73.154 100 90.962 81.376 85.623 82.574 71.750
200 88.715 87.165 79.348 75.344 59.673 200 88.194 74.738 84.246 77.869 58.904
300 88.604 71.054 86.236 74.168 53.010 300 87.954 70.676 83.750 71.874 51.768
400 88.199 67.091 85.579 71.479 48.825 400 87.222 67.051 82.483 68.803 48.243
500 87.724 64.172 84.897 68.671 46.777 500 87.356 63.717 81.543 66.169 46.229
602 87.203 83.937 66.774 61.168 45.600 592 87.133 80.339 64.528 61.134 45.139
Query VGG16 Query ViT
HSJA CGBA CGBA-H AHA TEA HSJA CGBA CGBA-H AHA TEA
100 93.966 92.154 85.805 84.770 75.634 100 80.454 76.885 73.074 65.501 65.347
200 89.884 90.899 81.599 78.445 61.148 200 79.981 74.702 69.511 58.863 54.851
300 89.665 73.174 90.516 74.359 54.977 300 79.606 53.951 74.139 63.655 49.164
400 88.943 68.538 89.880 70.559 50.940 400 79.599 49.857 72.839 61.300 46.139
500 88.919 65.569 89.563 67.745 48.943 500 78.761 46.572 71.291 58.598 44.347
588 88.389 88.362 66.578 62.126 47.484 554 79.626 69.954 57.710 44.494 43.422
ResNet50
ResNet101
VGG16
ViT
Figure 3: Average $\ell_{2}$ distance reduction across different architectures in a low-query regime. Higher values indicate improved performance.
Attack Success Rate. In Figures 4 and  5, we show how many images reach at least 50% and 75% distance reduction over queries respectively. A sharper increase in this fraction signifies that more pairs experience substantial improvement more quickly. Again, once the turning point is reached between a pair, CGBA-H is implemented for further refinement.
ResNet50
ResNet101
VGG16
ViT
Figure 4: Comparison of ASR of 50% distance reduction. Higher values indicate that a higher proportion of images reach a distance reduction of 50% sooner.
ResNet50
ResNet101
VGG16
ViT
Figure 5: Comparison of ASR of 75% distance reduction. Higher values indicate that a higher proportion of images reach a distance reduction of 75% sooner.
Success at a Fixed Low-Query Budget. Figure 6 reports, at a low-query budget of 500 queries, the proportion of image pairs that reach or exceed various distance-reduction thresholds. TEA maintains a consistently higher proportion of successful pairs across nearly all thresholds, suggesting that its early-stage perturbations secure significant distance reductions more rapidly.
ResNet50
ResNet101
VGG16
ViT
Figure 6: Cumulative distribution functions (CDFs) of distance reduction at 500 queries. Each curve represents the fraction of image pairs that achieve a given percentage reduction in the $\ell_{2}$ distance, with higher values indicating a more effective reduction method.
AUC Comparisons. We assess the efficiency of our aproach by computing the area under the median $\ell^{2}$-distance vs queries curve (AUC). Table 2 presents the AUC values for the low-query regime, up to the turning point. Across all architectures, TEA consistently yields lower AUC values.
Table 2: Average AUC values computed until the turning point across different architectures. Lower AUC values denote more effective early-stage distance reduction.
Model HSJA CGBA CGBA-H AHA TEA
ResNet50 53575.41 52130.74 45745.09 44332.55 35133.14
ResNet101 52269.70 49487.56 43865.80 43542.48 34189.94
VGG16 53053.06 53226.27 45261.02 44756.50 35790.23
ViT 44377.75 41097.13 36788.92 37356.25 30337.35
High-Query Performance with CGBA-H refinement. Our hybrid TEA+CGBA-H strategy consistently matches or surpasses the state-of-the-art. Table 3 lists the median $\ell_{2}$ distances upto 20000 queries. Note that CGBA-H is randomness-dependent. Further, it doesn’t re-explore when trapped in a narrow decision space and its stability can be unpredictable and subject to local decision boundary geometry. Figure 7 illustrates the median percentage reduction in the distance between the images.
Table 3: Median $\ell_{2}$ distances across different architectures. Here, TEA indicates using TEA until the turning point, and refining further with CGBA-H.
Query ResNet50 Query ResNet101
HSJA CGBA CGBA-H AHA TEA HSJA CGBA CGBA-H AHA TEA
1000 86.800 80.709 58.815 52.069 41.108 1000 86.162 76.005 55.824 51.642 40.090
2500 85.262 69.031 39.701 33.165 29.747 2500 85.314 63.477 38.043 33.259 29.965
5000 83.345 48.703 22.835 18.394 18.338 5000 84.202 41.437 22.470 18.780 18.932
10000 81.623 22.094 10.214 8.597 8.798 10000 82.443 17.623 10.341 9.021 9.864
15000 80.642 10.999 5.908 6.757 5.429 15000 81.873 8.936 6.341 6.937 6.281
20000 79.605 6.594 4.185 6.449 4.011 20000 80.886 5.689 4.507 6.644 4.560
Query VGG16 Query ViT
HSJA CGBA CGBA-H AHA TEA HSJA CGBA CGBA-H AHA TEA
1000 87.451 85.772 56.388 52.361 41.578 1000 79.212 64.516 49.394 49.541 38.482
2500 86.346 74.963 35.686 30.429 27.984 2500 78.739 44.789 32.573 34.762 28.134
5000 85.219 53.869 19.753 16.076 16.050 5000 78.170 25.118 18.566 22.197 18.507
10000 85.219 20.792 8.558 7.456 7.626 10000 77.786 10.319 9.299 11.267 9.416
15000 85.219 9.455 5.306 6.238 4.979 15000 77.087 5.920 5.987 8.180 6.301
20000 85.219 5.687 3.985 6.042 3.823 20000 76.727 4.258 4.635 7.404 4.955
ResNet50
ResNet101
VGG16
ViT
Figure 7: Comparison of median percentage decrease in $\ell^{2}$ distance across different architectures, with higher values indicating a more effective reduction method.
5 Conclusion
In this work, we introduced TEA, a targeted, hard-label, black-box adversarial attack that leverages edge information from a target image to efficiently produce adversarial examples perceptually closer to a source image in low-query settings. TEA initially employs a global search that preserves prominent edge structures across the target image, subsequently refining the perturbations via patch-wise updates. Empirical results demonstrate that TEA significantly reduces the $\ell_{2}$ distance between adversarial and source images, requiring over 70% fewer queries compared to current state-of-the-art methods. In addition, it also achieves reduced AUC and high ASR scores, indicating a consistently rapid reduction in distance to the source image across different models.
Limitations. (i) Randomness: Similar to state-of-the-art methods, TEA involves inherent randomness. In the patch-based refinement step, positions and sizes of patches are chosen randomly, leading to varying results across runs. Nevertheless, the overall performance remains stable when averaged over the thousand image pairs tested. (ii) Low-query focus: TEA is optimized specifically for scenarios with limited query budgets. Beyond this low-query regime, the method’s progress plateaus in a narrower decision space. However, the adversarial examples generated at this stage serve as strong initializations for subsequent refinement by other existing techniques.
Future Work. Several extensions could enhance TEA: (a) substituting edge information with other features such as textures, color distributions, or high-frequency components; (b) training a surrogate model based on query data already collected to guide subsequent perturbations and (c) developing defense mechanisms capable of mitigating attacks based on structure-preserving perturbations.
Code Availability
Our code is available at the following URL: https://github.com/mdppml/TEA.
Code of Ethics and Broader Impact Statement
We evaluate TEA exclusively on the publicly available ImageNet-1K validation set, which is distributed for non-commercial research and educational use under the ImageNet access agreement. The dataset contains no personally identifiable information, and we used it in accordance with the applicable license terms. As TEA reduces the number of required queries for targeted hard-label attacks in a low-query setting, it could be leveraged to more rapidly craft adversarial inputs against commercial or safety-critical vision systems. We encourage practitioners to adopt defense mechanisms that go beyond detecting frequency-based perturbations, explicitly incorporating checks for edge-informed distortions, and defenses attuned to other structural features, to guard against similar attacks.
Acknowledgements
This research was supported by the German Federal Ministry of Education and Research (BMBF) under project number 01ZZ2010 and partially funded through grant 01ZZ2316D (PrivateAIM). The authors acknowledge the usage of the Training Center for Machine Learning (TCML) cluster at the University of Tübingen.