### v3

# Accelerating Targeted Hard-Label Adversarial Attacks in Low-Query Black-Box Settings

Abstract
Deep neural networks for image classification remain vulnerable to adversarial examples — small, imperceptible perturbations that induce misclassifications. In black-box settings, where only the final prediction is accessible, crafting targeted attacks that aim to misclassify into a specific target class is particularly challenging due to narrow decision regions. Current state-of-the-art methods often exploit the geometric properties of the decision boundary separating a source image and a target image rather than incorporating information from the images themselves. In contrast, we propose Targeted Edge-informed Attack (TEA), a novel attack that utilizes edge information from the target image to carefully perturb it, thereby producing an adversarial image that is closer to the source image while still achieving the desired target classification. Our approach consistently outperforms current state-of-the-art methods across different models in low query settings (nearly 70% fewer queries are used), a scenario especially relevant in real-world applications with limited queries and black-box access. Furthermore, by efficiently generating a suitable adversarial example, TEA provides an improved target initialization for established geometry-based attacks.
I Introduction
Deep neural networks have achieved remarkable performance in image classification tasks, powering applications from autonomous systems [2, 6] to medical diagnostics [15, 20]. However, they have repeatedly been shown to be vulnerable to adversarial examples [19, 34]. These are small, often imperceptible perturbations to a correctly classified image that cause a misclassification. Although many of these attacks assume white-box access to a model’s internals, hard-label (decision-based) attacks offer a more challenging yet practical setting where only the top-1 predicted label is observed. This limited-feedback scenario commonly arises in commercial APIs [21]. In this realm, targeted attacks, which push the model’s prediction to a specific target class, are inherently more difficult since the decision regions corresponding to the specific target classes are usually narrower and more isolated.
In black-box settings, targeted hard-label attacks have become an active area of research, with several techniques proposed in the literature [10, 11, 36, 8, 9], with state-of-the-art methods relying on the geometry of the decision boundary separating the source image from the target class. Geometry-informed attacks traverse in a lower-dimensional space and fall into two categories: Boundary Tracing Attacks and Gradient Estimation Attacks. Boundary Tracing Attacks ([3, 4, 28, 33]) perform walks along the decision boundary while Gradient Estimation Attacks ([7, 24, 26, 27, 30, 35, 37]) perform the same walks but using information about the approximate tangent/normal to the decision boundary in a local neighborhood to their adversarial image. Although powerful for local refinement, both approaches tend to burn through queries when the source image lies far from a given adversarial image in a target class, wasting many queries before reaching a narrow region where local geometry can more effectively be leveraged.
Further, in practice, many real-world scenarios such as commercial pay-per-query APIs, impose severe constraints on the number of queries that can be made to a target model in a specified time. Often practical query limits may be on the order of a few hundred to fewer than a few thousand queries. Under these conditions, the limited feedback available makes it difficult to effectively use information about the local decision boundary geometry in the early stages of an attack. When the decision space is still wide, movement along restricted lower dimensional spaces leads to limited incremental progress, and gradient estimation methods often use queries in estimating local gradients by sampling predictions in a local neighborhood instead of moving. This raises a pressing need to develop methods that are efficient in the low-query regime. In a high query setting, this would also lead us to a good starting point to employ the geometry-informed methods since we arrive at a good adversarial point quickly and can use the geometry-informed information more effectively. To this end, we propose Targeted Edge-Informed Attack (TEA), a novel targeted adversarial attack designed for hard-label black-box scenarios within a restricted query budget. Rather than depending on the local properties of the decision boundary, TEA leverages the intrinsic features of the target image itself - specifically, the edge information obtained using Sobel filters [32] help identify prominent structural features in the target image. Edges encode high‑magnitude spatial gradients that delineate object boundaries [32]. Research has shown that early layers fire on oriented edge filters [1, 39], and that shape/edge cues remain predictive even when textures are suppressed [18]. The core idea is to preserve these low-level features, while applying perturbations to the non-edge regions of an image, allowing us to stay in the target class while pushing the adversarial image towards the source image. When our progress plateaus, evidenced by a series of consecutive queries that fail to achieve further reduction in distance while maintaining target class prediction, one can switch to current state-of-the-art geometry-informed methods for a local refinement procedure. Hence we make the following contributions:
TEA: We introduce an edge-informed perturbation strategy, enabling rapid progress toward the source image in the early stages of an adversarial attack when limited queries are available. TEA involves a two-step process: First, a global edge-informed search is performed, and then, edge-informed updates are applied to small patches using Gaussian weights.
Empirical validation under strict query budgets: We perform extensive evaluations on the ImageNet validation dataset [12] across four architectures (ResNet‑50 [20], ResNet‑101 [20], VGG16 [31], and ViT [13]) and an adversarially trained architecture (ResNet-50 [14]). Our attack consistently outperforms existing state‑of‑the‑art hard‑label methods, including HSJA, Adaptive History‑driven Attack (AHA) [25], CGBA, and CGBA‑H, under realistic query budgets (fewer than 1000 queries). To achieve a 60% reduction in distance from a target image to a source image, TEA required on average 251 queries across the four models - 70% fewer than AHA (the second fastest), which required 845 queries.
The rest of this paper is structured as follows. Section 2 reviews related work on targeted hard-label adversarial attacks. Section 3 describes our methodology, while Section 4 presents experimental results. Finally, Section 5 outlines directions for future research.
II Related Works
Early work in the realm of targeted hard-label adversarial attacks consisted of seminal work: Boundary Attack (BA) [3], which proposed a method of traversing the decision boundary that separates an adversarial image from a source image. Building on this framework, BiasedBA [4] incorporated directional priors, such as perceptual and low-frequency biases, to restrict the search to more promising regions. BA with Output Diversification Strategy (BAODS) [35] integrates diverse gradient-like signals into the exploration process of BA, thereby strengthening the original method.
Following these foundational methods, Hybrid Attack (HA) [33] employed a combination of heuristic search strategies to balance global exploration and local refinement. Advancing toward more gradient-centric techniques, qFool [26] leverages the local flatness of decision boundaries to streamline the attack process. Meanwhile, HopSkipJumpAttack (HSJA) [7] locally approximates the normal to the decision boundary to “jump off” from it before progressing toward the source image. Complementarily, Tangent Attack exploits locally estimated tangents of the decision surface to steer the adversarial perturbation toward the source image. Addressing the challenges posed by high-dimensional input spaces, Query Efficient Boundary Attack (QEBA) [24] projects gradient estimation into lower-dimensional subspaces, such as the frequency domain, thus significantly reducing the number of required queries.
Incorporating historical query information, Adaptive History-driven Attack (AHA) [25] adapts its search trajectory based on previous successes and failures, while Decision-based query Efficient Adversarial Attack based on boundary Learning (DEAL) [7] employs an evolutionary strategy that concentrates queries on promising regions of the input space. SurFree [28] demonstrated that using 2D planes and semicircular trajectories toward the source image was an effective strategy. This was based on the fact that under the assumption of a flat decision boundary, the point on the boundary that is closest to the source image is precisely where the semicircular trajectory intersects it. Building on this idea, Curvature-Aware Geometric black-box Attack (CGBA) and its variant that is more suited for targeted attacks - CGBA-H [30], use normal estimation at the decision boundary to select the traversal plane. To the best of our knowledge, CGBA-H serves as the current state of the art in decision-based targeted adversarial attacks.
III Methodology
In this section, we formalize the hard-label attack setting and detail our attack, which drastically reduces early query cost, especially when the target image is far from the source image, and exists in a wider decision space. Once the adversarial image achieves a good distance and reaches a narrow decision space, one can use the image as an initialization and continue refining with existing geometric-based methods.
Problem Statement. We consider a hard-label image classifier modeled by
$f:[0,1]^{C\times H\times W}\to\mathbb{R}^{K},$ (1)
where $C$ denotes the number of color channels, $H$ and $W$ are the image height and width, and the classifier distinguishes among $K$ classes. For any query image $x$, we do not observe its continuous output (e.g., logits or probabilities) but only the predicted label index
$\hat{y}(x)=\arg\max_{1\leq k\leq K}[f(x)]_{k}.$ (2)
Let $x_{s}$ be a source image correctly classified as $y_{s}$. In a targeted attack, we start with a target image $x_{t}$ (correctly classified as $y_{t}$). Our goal is to find an adversarial image $x_{\mathrm{adv}}$ that is as close as possible to $x_{s}$ (in the $\ell_{2}$ norm) while maintaining the target classification, i.e.,
$x_{\mathrm{adv}}=\arg\min_{x}\|x-x_{s}\|_{2},\quad\text{subject to}\quad\hat{y}(x)=y_{t}.$ (3)
We propose a two-part procedure for perturbing the target image $x_{t}$ to get closer to $x_{s}$ while maintaining adversarial requirements. First, a global edge-informed search coarsely aligns major image regions. Second, a patch-based edge-informed search further perturbs local regions in our adversarial image. Note that neither step leverages local decision boundary geometry or gradient estimation, which are typically beneficial only when the adversarial image lies near a narrow decision space and when movements towards the source image are unlikely to maintain classification. This approach avoids burning queries in a wider decision space.
III-A Global Edge-Informed Search
The global perturbation step aims to coarsely align $x_{t}$ towards $x_{s}$ by modifying predominantly smooth, non-edge areas, thereby preserving crucial edge structures that help maintain target classification.
Soft Edge Mask. To this end, we detect edges in $x_{t}$ using the Sobel operator [32], as depicted in Figure 1. A subsequent blurring operation yields a soft edge mask $M_{\mathrm{edge}}$, where $M_{\mathrm{edge}}(i,j)\in[0,1]$ has values close to 1 at edge locations and gradually transitions to 0 in smoother regions. We describe this in Algorithm 1.
Algorithm 1 Soft Edge Mask
1:  Inputs: Image $x$, edge thresholds $\{T_{\ell},T_{h}\}$, Gaussian blur kernel size $b$, intensity factor $\gamma$, small constant $\epsilon$
2:  Output: Soft edge mask $M_{\mathrm{edge}}$
3:  $x^{\text{gray}}\leftarrow\text{GrayScale}(x)$
4:  $(s_{x},s_{y})\leftarrow(\text{Sobel}(x^{\text{gray}},0),\;\text{Sobel}(x^{\text{gray}},1))$
5:  $G\leftarrow(s_{x}^{2}+s_{y}^{2})^{1/2}$
6:  $G\leftarrow 255\cdot\left(G/(max_{i,j}(G)+\epsilon)\right)$
7:  $\text{edge\_mask}(i,j)\leftarrow\begin{cases}255,&\text{if }T_{\ell}\leq G(i,j)\leq T_{h},\\ 0,&\text{otherwise}\end{cases}$
8:  $\text{blurred}\leftarrow\text{GaussianBlur}(\text{edge\_mask},(b,b))$
9:  $\text{norm}\leftarrow\text{Normalize}(\text{blurred})$
10:  $M_{\mathrm{edge}}\leftarrow\gamma\cdot\text{norm}$
11:  return $M_{\mathrm{edge}}$
Global Interpolation. Starting from $x_{0}=x_{t}$, we perform iterative updates
$x_{k+1}\leftarrow x_{k}+\alpha(x_{s}-x_{k})\odot\left(I-M_{\mathrm{edge}}\right),$ (4)
where $\odot$ denotes the element-wise (Hadamard) product. The scaling factor $\alpha$ is optimized via a momentum based search. Masking out the edge regions in this update ensures that the interpolation primarily affects smooth areas, thus preserving the overall structure necessary for target classification. The complete procedure is summarized in Algorithm 2. Once the improvements begin to stagnate, we transition to a local, patch-based refinement.
Algorithm 2 Global Edge-Informed Search
1:  Inputs: Source image $x_{s}$, target image $x_{t}$, soft edge mask $M_{\mathrm{edge}}$, tolerance $\tau$, maximum queries $qc_{\max}$, initial step factor $\eta$, momentum $\mu$
2:  Output: Adversarial image $x_{\mathrm{adv}}$ such that $\hat{y}(x_{\mathrm{adv}})=y_{t}$
3:  $x_{\text{current}}\leftarrow x_{t}$, $v\leftarrow 0$, and $d\leftarrow x_{s}-x_{t}$
4:  Set step size: $s\leftarrow\|d\|_{2}\cdot\eta$, and initialize query count $qc\leftarrow 0$
5:  while $qc<qc_{\max}$ and $\|s\|\geq\tau$
6:    $v\leftarrow\mu\cdot v+(1-\mu)\cdot d$
7:    $x_{\text{next}}\leftarrow x_{\text{current}}+s\cdot(v\odot(I-M_{\mathrm{edge}}))$
8:    $qc\leftarrow qc+1$
9:     if $\hat{y}(x_{\text{next}})=y_{t}$
10:       $x_{\text{current}}\leftarrow x_{\text{next}}$
11:       $s\leftarrow 1.1\cdot s$
12:    else
13:       break
14:  return $x_{\text{current}}$
III-B Patch-Based Edge-Informed Search
After the global interpolation, some subregions of the image may still display significant discrepancies from $x_{s}$. In the patch-based refinement, we then partition the image into randomly selected patches of random sizes
$\mathcal{P}\subseteq\{1,\dots,H\}\times\{1,\dots,W\},$ (5)
and for each patch, construct a local soft edge mask $M_{\mathcal{P}}$ by restricting $M_{\mathrm{edge}}$ to $\mathcal{P}$. Next, starting from our adversarial image $x_{k}$, we apply a local interpolation
$\widetilde{x}(\beta)=x_{k}+\beta(x_{s}-x_{k})\odot G\odot(I-M_{\mathcal{P}}),$ (6)
and search for the largest $\beta$ such that the classifier still predicts the target label, $\hat{y}(\widetilde{x}(\beta))=y_{t}$. Here, $G$ is a Gaussian weighting function over the patch that smoothly downweights updates near patch borders, helping to avoid artificial edges introduced by patch boundaries. We repeat this for different patches until a termination criterion is met (e.g., 25 consecutive iterations with no further improvement). The corresponding illustration is depicted in Figure 1 and the relevant pseudocode is provided in Algorithm 3.
Figure 1: Overview of Patch-Based Edge-Informed Search. Edge information from the target image, obtained via the Sobel operator, is first blurred to generate a soft edge mask. A square patch is then selected and a Gaussian weighting function is applied. In the bottom right panel, the intensity of the modification is illustrated: dark red regions remain largely unchanged, while light green regions receive a more pronounced update. The lack of changes near the patch borders helps prevent the introduction of artificial edges.
Figure 2 offers a visual overview of TEA. The figure depicts the adversarial image throughout our method, and displays a hotspot of individual pixel differences from the source image.
Figure 2: Visualization of TEA on a source–target image pair. The target image (initially classified as Bee-eater) is perturbed to resemble the source image (classified as Spoonbill), while preserving its original Bee-eater label. Global Edge-Informed Search efficiently applies edge-aware perturbations using only $15$ queries to achieve a $\approx$20% reduction in distance to the source. Patch-Based Edge-Informed Search introduces localized, edge-aware modifications to small image regions, as seen in the hotspot of changes. Further refinement utilizing CGBA-H is illustrated in the narrow decision space.
Algorithm 3 Patch-Based Edge-Informed Search
1:  Inputs: Source image $x_{s}$, current adversarial image $x_{\mathrm{adv}}$, soft edge mask $M_{\mathrm{edge}}$ (from create_soft_edge_mask), minimum patch size $p_{\min}$, maximum patch size $p_{\max}$, step factor $\eta$, momentum $\mu$, maximum calls $N_{\max}$
2:  Output: Refined adversarial image $x_{\mathrm{adv}}$
3:  $n_{\text{break}}\leftarrow 0$
4:  while $n_{\text{break}}<25$
5:     $D\leftarrow\text{AvgPool}\left(|x_{s}-x_{\mathrm{adv}}|\right)$
6:     Select high-difference indices from $D$ and choose a random center $(i_{c},j_{c})$
7:     $p\leftarrow\mathrm{randInt}(p_{\min},\,p_{\max})$;   $\mathcal{P}\leftarrow\{(i,j)\mid|i-i_{c}|\leq\lfloor p/2\rfloor,\;|j-j_{c}|\leq\lfloor p/2\rfloor\}$
8:     $M_{\mathcal{P}}(i,j)\leftarrow\begin{cases}1,&(i,j)\in\mathcal{P},\\ 0,&\text{otherwise.}\end{cases}$
9:     $x_{\text{patch}}\leftarrow\mathrm{Patch}(x_{\mathrm{adv}},\mathcal{P})$
10:     $m_{\text{patch}}\leftarrow 0,\quad d_{\text{base}}\leftarrow\|x_{s}-x_{\mathrm{adv}}\|_{2}$
11:     for iteration =1=1 to $N_{\max}$
12:        $d_{\text{local}}\leftarrow\mathrm{Patch}(x_{s},\mathcal{P})-x_{\text{patch}}$
13:        $m_{\text{patch}}\leftarrow\mu\cdot m_{\text{patch}}+(1-\mu)\cdot d_{\text{local}}$
14:        $s_{\text{patch}}\leftarrow\eta\cdot\|x_{s}-x_{\mathrm{adv}}\|_{2}$
15:        $G\leftarrow\mathrm{GaussianWeight}((i_{c},j_{c}),\sigma=p/3)$
16:        $\Delta x\leftarrow s_{\text{patch}}\cdot m_{\text{patch}}\odot\left(G\odot\left(1-(M_{\mathrm{edge}}\odot M_{\mathcal{P}})\right)\right)$
17:        $x_{\text{patch}}^{+}\leftarrow x_{\text{patch}}+\Delta x$
18:        $x_{\mathrm{temp}}\leftarrow x_{\mathrm{adv}}-x_{\mathrm{adv}}\odot M_{\mathcal{P}}+x_{\text{patch}}^{+}\odot M_{\mathcal{P}}$
19:        $d_{\text{new}}\leftarrow\|x_{s}-x_{\mathrm{temp}}\|_{2}$
20:        if $d_{\text{new}}\geq 0.999\cdot d_{\text{base}}$
21:           break
22:        if $\hat{y}(x_{\mathrm{temp}})=y_{t}$
23:           $x_{\text{patch}}\leftarrow x_{\text{patch}}^{+}$
24:           $x_{\mathrm{adv}}\leftarrow\mathrm{Replace}(x_{\mathrm{adv}},\mathcal{P},x_{\text{patch}})$
25:           $d_{\text{base}}\leftarrow d_{\text{new}}$
26:           $n_{\text{break}}\leftarrow 0$
27:        else
28:           $n_{\text{break}}\leftarrow n_{\text{break}}+1$
29:           break
30:  return $x_{\mathrm{adv}}$
IV Experiments
In this section, we present our empirical evaluation benchmarking TEA against existing targeted hard-label attacks. In what follows, we describe our experimental setup and metrics, then detail the performance of each method under a range of query budgets. Our findings indicate that TEA achieves efficient distance reduction to the source image, particularly in the early query regime, before switching to the current state-of-the-art geometry-based attack CGBA-H for further refinement in a narrow decision space.
IV-A Setup and Metrics
Computational Resources. The experiments were executed on a High-performance computing (HPC) Cluster. Each node on the cluster consisted of four NVIDIA GeForce GTX 1080 Ti GPUs (one GPU was allocated per source-target pair for a given attack).
Dataset and Image Pairs. We randomly sample 1000 source-target image pairs from the ImageNet ILSVRC2012 validation set, ensuring each pair contains images from distinct classes. All images are resized to $3\times 224\times 224$. Each pair $(x_{s},x_{t})$ contains a source image $x_{s}$, correctly classified under its label, and a target image $x_{t}$, also correctly classified under a different label. The goal is to modify $x_{t}$ to approach $x_{s}$ under the $\ell_{2}$ norm while keeping the prediction unchanged.
Target Models. We evaluate our approach on four well-known classifiers: ResNet50 and ResNet101 (CNNs of varying depth), VGG16 (a CNN composed of repeated convolutional blocks), and ViT (a vision transformer that processes images as sequences of patches). These models represent a diverse set of architectures.
Compared Methods. We compare our method to four targeted hard-label attacks: HSJA, AHA, CGBA, and CGBA-H.For each method, the $\ell_{2}$ distance to the source image is recorded after each set of queries, and along with the classification label of the perturbed target image.
Evaluation Metrics. Performance is quantified using three metrics. First, we compute the $\ell_{2}$ distance from $x_{s}$ to the adversarial example generated for each of the 1000 image pairs as queries progress. Second, the attack success rate (ASR) is defined as the fraction of image pairs for which an $\alpha\%$ reduction in $\ell_{2}$ distance between adversarial image and source image is achieved as compared to the target image and source image. Third, we also measure the performance of each method when there is a set fixed low-query budget by measuring the ASRs for all $\alpha$ when each method has used up $500$ queries. Finally, we integrate the $\ell_{2}$ distance versus queries curve to obtain the area under the curve (AUC), which provides an aggregated measure of how rapidly the distance is reduced.
Implementation Details. Our implementation employs a two-stage process. In the first stage, we perform our proposed methodology TEA, where edge-aware distortions are applied to the target image while preserving its target classification, allowing a rapid reduction in distance to the source image. In the second stage, once the perturbations have brought the image into a narrow decision space, the method switches to CGBA-H for further refinement (denoted as TEA throughout the study). The last query performed with TEA is referred to as the turning point throughout the study. We switch to CGBA-H since it consistently performs better than other methods in a high query setting across different architectures.
IV-B Results
TABLE I: Median $\ell_{2}$ distances computed until the turning point across different architectures.
Lower $\ell_{2}$ values denote faster adversarial example generation.
Query ResNet50 Query ResNet101
HSJA CGBA CGBA-H AHA TEA HSJA CGBA CGBA-H AHA TEA
100 92.797 88.857 83.104 80.438 75.2155 100 90.962 85.623 81.376 82.574 74.5036
200 88.715 87.165 79.348 75.344 62.0338 200 88.194 84.246 74.738 77.869 61.3794
300 88.604 86.236 74.168 71.054 55.5589 300 87.954 83.750 71.874 70.676 54.9321
400 88.199 85.579 71.479 67.091 51.9530 400 87.356 82.483 68.803 67.051 51.3780
500 87.724 84.897 68.671 64.172 49.8352 500 87.222 81.543 66.169 63.717 49.3863
602 87.203 83.937 66.774 61.168 45.6000 592 87.133 80.339 64.528 61.134 45.1390
Query VGG16 Query ViT
HSJA CGBA CGBA-H AHA TEA HSJA CGBA CGBA-H AHA TEA
100 93.966 92.154 85.805 84.770 77.2608 100 80.454 76.885 73.074 68.863 67.5654
200 89.884 90.899 81.599 78.445 63.6530 200 79.981 74.702 69.511 65.501 56.1255
300 89.665 90.516 74.359 73.174 56.9954 300 79.606 74.139 63.655 63.951 50.8059
400 88.943 89.880 70.559 68.538 53.4018 400 79.599 72.839 61.300 59.857 47.8666
500 88.919 89.563 67.745 65.569 51.3763 500 78.761 71.291 58.598 56.572 46.2528
588 88.389 88.362 66.578 62.126 47.4840 554 78.626 69.954 57.710 54.494 43.4220
ResNet50
ResNet101
VGG16
ViT
Figure 3: Average $\ell_{2}$ distance reduction across different architectures in a low-query regime. Higher values indicate improved performance.
Average $\ell_{2}$-Distance vs. Queries. Table I presents a comparative analysis of the median $\ell_{2}$ distances achieved in the low-query regime—specifically, within the range defined by the average turning point computed over one thousand image pairs. This analysis underscores the rapid decrease in perturbation norm facilitated by TEA during the early query stages. Figure 3 extends this evaluation by depicting the average percentage reduction in $\ell_{2}$ distance against queries used. In the plots, the initial inflection point marks TEA’s transition from global exploration to patch-based refinement.
ResNet50
ResNet101
VGG16
ViT
Figure 4: Comparison of ASR of 50% distance reduction. Higher values indicate that a higher proportion of images reach a distance reduction of 50% sooner.
ResNet50
ResNet101
VGG16
ViT
Figure 5: Comparison of ASR of 75% distance reduction. Higher values indicate that a higher proportion of images reach a distance reduction of 75% sooner.
Attack Success Rate. In Figures 4 and  5, we show how many images reach at least 50% and 75% distance reduction over queries respectively. A sharper increase in this fraction signifies that more pairs experience substantial improvement more quickly. Again, once the turning point is reached between a pair, CGBA-H is implemented for further refinement.
Success at a Fixed Low-Query Budget. Figure 6 reports, at a low-query budget of 500 queries, the proportion of image pairs that reach or exceed various distance-reduction thresholds. TEA maintains a consistently higher proportion of successful pairs across nearly all thresholds, suggesting that its early-stage perturbations secure significant distance reductions more rapidly.
ResNet50
ResNet101
VGG16
ViT
Figure 6: Cumulative distribution functions (CDFs) of distance reduction at 500 queries. Each curve represents the fraction of image pairs that achieve a given percentage reduction in the $\ell_{2}$ distance, with higher values indicating a more effective reduction method.
AUC Comparisons. We assess the efficiency of our aproach by computing the area under the median $\ell_{2}$-distance vs queries curve (AUC). Table II presents the AUC values for the low-query regime, up to the turning point. Across all architectures, TEA consistently yields lower AUC values.
TABLE II: Average AUC values computed until the turning point across different architectures. Lower AUC values denote more effective early-stage distance reduction.
Model HSJA CGBA CGBA-H AHA TEA
ResNet50 53575.41 52130.74 45745.09 44332.55 35133.14
ResNet101 52269.70 49487.56 43865.80 43542.48 34189.94
VGG16 53053.06 53226.27 45261.02 44756.50 35790.23
ViT 44377.75 41097.13 36788.92 37356.25 30337.35
Evaluation on an Adversarially Trained Model. In Figure 7, we present a comparative analyses of performance against an adversarial trained model (Resnet50 from MardyLab [14]) in the low-query region. We see that TEA consistently achieves more distortion, while also noting that all methods perform significantly worse compared to the standard Resnet50 model as seen in Figure 3.
Refer to caption Refer to caption
Figure 7: Comparison on an adversarially trained model. Left: median percentage decrease in $\ell_{2}$ distance against number of queries, with higher values indicating a more effective reduction method. Right: median $\ell_{2}$ distance against number of queries, with lower values indicating a more effective reduction method.
Edge Ablation Study To better understand the role of edge preservation in our method, we perform an ablation study by comparing three variants of TEA under identical settings. TEA perturbs non-edge pixels, preserving edge regions with a soft edge mask. INV-TEA inverts TEA’s soft edge mask to perturb primarily edge-like pixels. HALF-TEA perturbs non-edge and edge regions equally, using the same editable-pixel budget as TEA, and serves as a simple control baseline. Table III summarizes results in the low-query regime: median $\ell_{2}$ distance, ASR at $50\%$ distance reduction, and ASR at $75\%$ distance reduction. Across all four architectures, TEA achieves the lowest median $\ell_{2}$ and the highest success rates overall; HALF-TEA is consistently in between, and INV-TEA performs the worst, indicating that prioritizing non-edge regions proves crucial for early attack progress while preserving the semantic structure.
TABLE III: Edge ablation in the low-query regime.
Per architecture and budget we report median $\ell_{2}$ (lower is better), ASR at 50%, and ASR at 75% (higher is better).
Query ResNet-50 Query ResNet-101
$\ell_{2}$ ASR@50% ASR@75% $\ell_{2}$ ASR@50% ASR@75%
TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv
100 75.1377 87.3465 87.5367 44.0 23.6 22.8 2.0 0.2 1.3 100 74.4637 86.1526 86.8610 44.4 24.1 23.2 2.8 0.3 1.0
200 61.8307 73.1697 73.3559 70.5 49.7 45.2 6.5 0.7 2.6 200 61.3517 71.7322 72.6071 72.4 53.3 47.3 6.9 0.9 3.1
300 55.3574 65.0474 66.5448 81.8 67.8 61.8 10.7 3.3 4.4 300 54.9709 63.5007 65.8087 83.9 69.7 62.0 11.8 3.2 4.7
400 51.8838 60.0257 63.2027 86.0 77.6 69.0 15.6 5.9 6.1 400 51.3084 58.5084 62.4517 87.5 77.6 69.0 17.1 6.7 6.0
500 49.8877 56.8181 61.6506 87.5 81.8 71.3 19.2 9.5 6.9 500 49.3153 55.3806 60.8018 89.6 82.6 72.2 19.7 9.5 6.5
Query VGG16 Query ViT
$\ell_{2}$ ASR@50% ASR@75% $\ell_{2}$ ASR@50% ASR@75%
TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv
100 77.4206 90.7514 89.8802 38.1 17.6 20.8 2.0 0.3 0.8 100 67.4874 76.4322 77.6473 57.8 43.4 40.4 5.2 0.9 2.8
200 63.7679 75.3909 75.5661 66.3 45.0 40.8 5.8 0.7 2.3 200 56.2801 63.8529 66.0835 80.0 69.8 60.8 11.8 3.6 6.7
300 57.0986 66.5727 68.5222 78.7 64.8 56.9 10.8 2.9 3.9 300 50.8057 57.2091 60.8559 88.1 82.8 71.0 17.5 7.7 9.7
400 53.4352 61.0717 65.1434 83.3 74.9 62.9 16.8 7.4 5.0 400 47.9956 53.3073 58.3993 90.6 88.1 74.5 22.9 12.7 12.1
500 51.4759 57.5518 63.4061 84.7 79.2 66.6 19.5 11.4 6.8 500 46.3913 51.0218 57.2895 92.1 90.6 76.0 26.1 16.2 13.0
Randomness and Stability Across Patch Selection Our patch-based refinement uses random patch locations and sizes, introducing stochasticity. To quantify this effect, we executed five independent runs per model. For each run and query budget, we computed the average $\ell_{2}$ distance across all pairs; we then summarized the five runs by reporting mean $\pm$ standard deviation (std) of the average $\ell_{2}$ distance across all pairs. As seen in Table IV, across all models and budgets, variability is small relative to the mean. The standard deviation is typically below $0.3$ (well under $1\%$ of the mean).
TABLE IV: Randomness analysis across five runs: $\ell_{2}$ (mean $\pm$ std).
Budget ResNet50 ResNet101 VGG16 ViT
50 88.1415 $\pm$ 0.0254 87.0992 $\pm$ 0.0637 90.6675 $\pm$ 0.0553 78.7152 $\pm$ 0.0277
100 75.2155 $\pm$ 0.2338 74.5036 $\pm$ 0.0676 77.2608 $\pm$ 0.1085 67.5654 $\pm$ 0.0652
150 67.3196 $\pm$ 0.2301 66.6368 $\pm$ 0.1022 69.1007 $\pm$ 0.1479 60.6342 $\pm$ 0.1187
200 62.0338 $\pm$ 0.2607 61.3794 $\pm$ 0.1318 63.6530 $\pm$ 0.1896 56.1255 $\pm$ 0.0829
250 58.2941 $\pm$ 0.2458 57.6515 $\pm$ 0.1356 59.7896 $\pm$ 0.2565 53.0414 $\pm$ 0.0537
300 55.5589 $\pm$ 0.1995 54.9321 $\pm$ 0.1397 56.9954 $\pm$ 0.2654 50.8059 $\pm$ 0.0627
400 51.9530 $\pm$ 0.1907 51.3780 $\pm$ 0.1564 53.4018 $\pm$ 0.2513 47.8666 $\pm$ 0.0654
500 49.8352 $\pm$ 0.1522 49.3863 $\pm$ 0.1306 51.3763 $\pm$ 0.2361 46.2528 $\pm$ 0.0899
Lower Resolution Performance We evaluate TEA, CGBA, CGBA-H and AHA at two lower input sizes, $128\times 128$ and $64\times 64$. The Tables V and VI report the median $\ell_{2}$ distance to the source image achieved at fixed query budgets; lower is better. VGG-16 and ViT required standard architectural adjustments at reduced dimensions.
TABLE V: Median $\ell_{2}$ distances at $\times128$.
Query ResNet50 Query ResNet101
AHA CGBA CGBA-H TEA AHA CGBA CGBA-H TEA
50 52.812610 56.892204 56.892204 53.891792 50 52.325765 55.829930 55.830184 52.836270
100 49.792591 52.638543 52.610928 44.656296 100 49.256166 51.690833 52.297024 43.908848
150 47.580821 48.751302 48.900604 40.200085 150 47.165379 47.665892 48.747030 39.356266
200 45.816245 47.670539 47.708520 37.588080 200 45.418667 46.499661 48.731472 36.744068
250 44.242092 47.291906 47.416906 35.637354 250 43.904684 45.875023 48.286582 34.803333
300 42.705044 45.310037 45.394565 34.083014 300 43.015997 43.947135 46.251110 33.364790
350 41.704023 45.917060 46.053402 32.915845 350 42.285534 45.637712 48.112825 32.363468
400 40.606535 43.876574 44.083020 31.711626 400 41.279571 43.208419 47.001603 31.200894
Query VGG16 Query ViT
AHA CGBA CGBA-H TEA AHA CGBA CGBA-H TEA
50 53.742676 57.595045 57.595044 54.653168 50 49.717520 61.999604 61.999605 51.518260
100 50.687319 53.132007 53.124912 45.101308 100 46.560914 56.845693 56.760291 42.468795
150 48.452131 48.769476 48.688020 40.641963 150 44.133746 51.938031 51.849316 37.765469
200 46.376375 47.648656 47.548296 37.729822 200 42.135314 52.195369 52.190025 34.882641
250 44.719990 46.665404 46.692002 35.775914 250 40.692779 50.410471 50.458582 32.847132
300 43.824315 44.433300 44.557708 34.574623 300 39.469095 47.456320 47.224463 31.225525
350 42.589694 45.370771 45.405220 33.251792 350 38.537397 49.233241 49.101166 30.127403
400 41.597263 43.446274 43.461510 32.151386 400 37.460420 47.532119 47.482755 29.117229
TABLE VI: Median $\ell_{2}$ distances at $\times64$.
Query ResNet50 Query ResNet101
AHA CGBA CGBA-H TEA AHA CGBA CGBA-H TEA
50 28.562640 31.263057 31.263184 29.709433 50 28.527201 31.240249 31.240327 29.828940
100 26.583674 28.666067 28.734703 25.884418 100 26.545651 28.491136 28.543173 26.137666
150 25.094041 26.041276 26.107465 23.768843 150 25.101155 25.824065 25.971276 23.934921
200 23.449901 24.898503 25.231694 21.971898 200 23.850520 24.798062 25.058821 22.132178
250 22.191710 24.117243 24.478960 20.671545 250 22.805084 23.948980 24.090509 20.863024
Query VGG16 Query ViT
AHA CGBA CGBA-H TEA AHA CGBA CGBA-H TEA
50 28.318550 30.844857 30.837613 29.530338 50 30.081730 32.758422 32.758422 31.046581
100 26.390916 28.302722 28.346767 25.606639 100 27.818672 29.806229 29.807765 27.077744
150 24.731502 25.578313 25.651390 23.405260 150 26.072833 26.316862 26.340219 24.609781
200 23.409042 24.719593 24.810609 21.847981 200 24.510567 24.761248 24.786895 22.262164
250 22.453166 23.539063 23.687779 20.527478 250 23.248782 23.717796 23.829428 20.461638
As observable in the results, at lower input resolution sizes, each patch modification distorts a larger fraction of the target image and thus contributes more to the overall $\ell_{2}$-norm reduction. Consequently, maintaining the adversariality of the image becomes much harder, and TEA therefore terminates earlier than for higher resolutions.
Perceptual Metrics While adversarial examples are generated to trick ML models, it is worth considering their impact on human perception [5, 17]. We report SSIM [38] and FSIM [40] alongside distortion at a fixed budget of 400 queries. As shown in Table VII, TEA achieves high distortion while slightly improving FSIM. A small drop in SSIM is expected, since TEA’s perturbations are more structured. Prior image quality assessment (IQA) studies [40, 16] indicate FSIM correlates more strongly with human perception than SSIM, which helps contextualise these differences.
TABLE VII: SSIM and FSIM at 400 queries on 1000 source—target pairs.
Lower $\ell_{2}$ indicates smaller perturbations; higher SSIM/FSIM indicates greater perceptual similarity to the source image.
ResNet50 ResNet101
Metric CGBA CGBA-H AHA TEA CGBA CGBA-H AHA TEA
$\ell_{2}$ 85.579 71.479 64.172 51.9530 82.483 68.803 67.051 51.3780
SSIM 0.531765 0.530459 0.231784 0.478806 0.525847 0.528782 0.223511 0.491950
FSIM 0.251337 0.252585 0.232117 0.294672 0.252746 0.251235 0.223831 0.288153
VGG16 ViT
Metric CGBA CGBA-H AHA TEA CGBA CGBA-H AHA TEA
$\ell_{2}$ 89.880 70.559 68.538 53.4018 72.839 61.300 59.857 47.8666
SSIM 0.534783 0.532189 0.231784 0.457879 0.580527 0.582544 0.201767 0.548110
FSIM 0.251129 0.251239 0.232117 0.305919 0.220927 0.220213 0.202171 0.257571
High‐Query Performance with CGBA-H refinement. Our hybrid TEA+CGBA‑H strategy consistently matches or surpasses the state-of-the-art. Table VIII lists the median $\ell_{2}$ distances upto 20000 queries. Note that since CGBA-H is randomness-dependent, it doesn’t re-explore when trapped in a narrow decision space and its stability is subject to local decision boundary geometry. Figure 8 illustrates the median percentage reduction in the distance between the images.
TABLE VIII: Median $\ell_{2}$ distances across different architectures.
Here, TEA indicates using TEA until the turning point, and refining further with CGBA-H.
Query ResNet50 Query ResNet101
HSJA CGBA CGBA-H AHA TEA HSJA CGBA CGBA-H AHA TEA
1000 86.800 80.709 58.815 52.069 41.108 1000 86.162 76.005 55.824 51.642 40.090
2500 85.262 69.031 39.701 33.165 29.747 2500 85.314 63.477 38.043 33.259 29.965
5000 83.345 48.703 22.835 18.394 18.338 5000 84.202 41.437 22.470 18.780 18.932
10000 81.623 22.094 10.214 8.597 8.798 10000 82.443 17.623 10.341 9.021 9.864
15000 80.642 10.999 5.908 6.757 5.429 15000 81.873 8.936 6.341 6.937 6.281
20000 79.605 6.594 4.185 6.449 4.011 20000 80.886 5.689 4.507 6.644 4.560
Query VGG16 Query ViT
HSJA CGBA CGBA-H AHA TEA HSJA CGBA CGBA-H AHA TEA
1000 87.451 85.772 56.388 52.361 41.578 1000 79.212 64.516 49.394 49.541 38.482
2500 86.346 74.963 35.686 30.429 27.984 2500 78.739 44.789 32.573 34.762 28.134
5000 85.219 53.869 19.753 16.076 16.050 5000 78.170 25.118 18.566 22.197 18.507
10000 85.219 20.792 8.558 7.456 7.626 10000 77.786 10.319 9.299 11.267 9.416
15000 85.219 9.455 5.306 6.238 4.979 15000 77.087 5.920 5.987 8.180 6.301
20000 85.219 5.687 3.985 6.042 3.823 20000 76.727 4.258 4.635 7.404 4.955
ResNet50
ResNet101
VGG16
ViT
Figure 8: Comparison of median percentage decrease in $\ell_{2}$ distance across different architectures — higher values indicate a more effective reduction method.
Source-Target Similarity Analysis. To examine how the performance of TEA depends on the similarity between source and target images, we group the $1000$ source-image pairs by structural similarity. For each pair $(x_{s},x_{t})$ we compute the SSIM score and sort all pairs by SSIM, partitioning them into ten equally sized bins. Within each bin, and model, we measure the percentage $\ell_{2}$ reduction at a fixed budget of $400$ queries. We then average this quantity over all pairs in the same bin and plot the resulting curves (Figure 9). We see that all methods have relatively identical performance variability, with improved performance across structurally similar source-target image pairs. This indicates that structural similarity does not favor TEA over the other methods.
ResNet50
ResNet101
VGG16
ViT
Figure 9: Effect of source–target structural similarity on attack performance. We group image pairs into ten bins by SSIM and report the average percentage decrease in $\ell_{2}$ distance after 400 queries. All methods benefit similarly from higher structural similarity.
Edge–Density Analysis. To assess how TEA behaves under different levels of structural detail in the source and target images, we perform an edge–density based stratification of the source–target pairs. For each image, we convert it to grayscale, compute Sobel gradients, and form the normalized gradient magnitude $\tilde{g}(p)\in[0,1]$ for each pixel $p$. Pixels with $\tilde{g}(p)>0.2$ are treated as edge pixels, and the edge density $d(x)$ is defined as the fraction of pixels classified as edges. We collect $d(x_{s})$ and $d(x_{t})$ for all source-target image pairs and set a global dense/sparse threshold $\tau_{\text{dense}}$ to be the median density over all $2000$ images, yielding $\tau_{\text{dense}}\approx 0.110242$ in our case. Images with $d(x)>\tau_{\text{dense}}$ are labeled dense, and the rest sparse. Each pair $(x_{s},x_{t})$ is thus assigned to one of four edge–pattern regimes: dense–sparse, dense–dense, sparse–sparse, or sparse–dense. For each regime, and model, we report the average percentage $\ell_{2}$ reduction (in %) at a fixed budget of $400$ queries. The resulting curves in Figure 10 show that performance variability is similar across methods, with higher success when moving towards a sparse source image than a dense one, and that TEA maintains a consistent advantage across all regimes.
ResNet50
ResNet101
VGG16
ViT
Figure 10: Effect of source–target edge density on attack performance. Each image is classified as dense or sparse based on a Sobel edge–density threshold, inducing four edge–pattern regimes (dense–dense, dense–sparse, sparse–dense, sparse–sparse). For each regime, and model, we report the average percentage decrease in $\ell_{2}$ distance after 400 queries, illustrating how performance varies with the edge richness of the source–target pairs.
Zero-Shot CLIP. To assess whether TEA also transfers to more modern vision-language models, we additionally evaluate it against a zero-shot CLIP classifier on the ImageNet validation set. We use the ViT-B/32 variant of CLIP [29, 13] in its standard zero-shot configuration. Following the usual protocol, we construct a linear zero-shot head by encoding multiple natural-language templates for each ImageNet class (e.g., “a photo of a {}.”, “a close-up photo of a {}.”) with the CLIP text encoder, normalizing and averaging the resulting embeddings to obtain a single prototype per class. Stacking these prototypes yields a fixed weight matrix, and logits are obtained by taking scaled dot products between normalized image features and these class prototypes. Our implementation relies on the official CLIP PyTorch code-base and its recommended preprocessing.
In Figure 11, we show the average percentage decrease in $\ell_{2}$ distance and the average $\ell_{2}$ distance as a function of the query budget. TEA retains a clear advantage, consistently achieving larger perturbation reductions than the baselines within the same limited query budget, indicating that its strategy remains effective even for modern zero-shot vision-language models.
Refer to caption Refer to caption
Figure 11: Comparison on a zero-shot CLIP (ViT-B/32) classifier on ImageNet. Left: average percentage decrease in $\ell_{2}$ distance against number of queries, with higher values indicating a more effective reduction method. Right: average $\ell_{2}$ distance against number of queries, with lower values indicating a more effective reduction method.
Other Datasets. To further assess the robustness and generality of TEA beyond ImageNet, we additionally evaluate all tested methods on two standard classification benchmark datasets: CIFAR-100 [23] and the Intel Image Classification dataset [22]. CIFAR-100 contains 100 object categories with lower-resolution images while the Intel dataset consists of six scene categories (e.g., buildings, forest, sea). In both cases, we again attack the four models, ResNet-50, ResNet-101, VGG-16, and ViT, and measure performance in terms of the relative $\ell_{2}$-distance reduction between the source and target images. Figures 12 and 13 report the average $\ell_{2}$-distance reduction at a fixed low-query budget of $400$ across all four models. Consistent with our results on the ImageNet dataset, TEA achieves larger perturbation reductions on both datasets, indicating that its edge-aware strategy remains effective across diverse data distributions and resolutions.
ResNet50
ResNet101
VGG16
ViT
Figure 12: Average $\ell_{2}$ distance reduction across different architectures in a low-query regime on the CIFAR-100 dataset.
ResNet50
ResNet101
VGG16
ViT
Figure 13: Average $\ell_{2}$ distance reduction across different architectures in a low-query regime on the Intel Image Classification dataset.
V Conclusion
In this work, we introduced TEA, a targeted, hard-label, black-box adversarial attack that leverages edge information from a target image to efficiently produce adversarial examples perceptually closer to a source image in low-query settings. TEA initially employs a global search that preserves prominent edge structures across the target image, subsequently refining the perturbations via patch-wise updates. Empirical results demonstrate that TEA significantly reduces the $\ell_{2}$ distance between adversarial and source images, requiring over 70% fewer queries compared to current state-of-the-art methods. In addition, it also achieves reduced AUC and high ASR scores, indicating a consistently rapid reduction in distance to the source image across different models.
Limitations. TEA is specifically designed for the targeted, hard-label, low-query black-box setting, and its main advantage lies in rapidly improving performance under tight query budgets. When TEA is used merely as an initialization and followed by stronger high-query attacks in regimes where many queries are available, we observe that the final performance is comparable to that of the underlying state-of-the-art methods, rather than strictly better. In addition, while TEA preserves global edge structure and achieves low $\ell_{2}$ distances, the patch-wise updates can still introduce localized artifacts that are visible to human observers. This limitation is shared with other norm-bounded attacks in general, but in TEA the patch-based nature of the perturbations can make these local changes particularly noticeable.
Future Work. Several extensions could enhance TEA: (a) substituting edge information with other features such as textures, color distributions, or high-frequency components; (b) training a surrogate model based on query data already collected to guide subsequent perturbations and (c) developing defense mechanisms capable of mitigating attacks based on structure-preserving perturbations.
Code Availability
Our code is available at the following URL: https://github.com/mdppml/TEA.
Code of Ethics and Broader Impact Statement
We evaluate TEA exclusively on the publicly available ImageNet-1K validation set, which is distributed for non-commercial research and educational use under the ImageNet access agreement. The dataset contains no personally identifiable information, and was used in accordance with the license terms. As TEA reduces the number of required queries for targeted hard-label attacks in a low-query setting, it could be leveraged to more rapidly craft adversarial inputs against commercial or safety-critical vision systems. We encourage practitioners to adopt defense mechanisms that go beyond detecting frequency-based perturbations, explicitly incorporating checks for edge-informed distortions, and defenses attuned to other structural features, to guard against similar attacks.
LLM Usage Considerations
LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality. All technical ideas, experiments, analyses, and citations were conceived, implemented, and validated by the authors.
Acknowledgments
This research was supported by the German Federal Ministry of Education and Research (BMBF) under project number 01ZZ2010 and partially funded through grant 01ZZ2316D (PrivateAIM). The authors acknowledge the usage of the Training Center for Machine Learning (TCML) cluster at the University of Tübingen.